Express the number appearing in the following statements in standard form.
 1 micron is equal to 1/1000000 m

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1 micron = 1/1000000
= 1/10^6
= 1×10^-6

Thanks